5.A balloon has a volume of 2.25 L when the temperature is 400C. What will be the volume if the balloon is cooled to 120C? Assume the pressure to be constant.


1
Expert's answer
2022-05-18T02:09:14-0400

Given:

V1 = 2.25 L

T1 = 40°C + 273.15 = 313.15 K

P = constant

V2 = unknown

T2 = 12°C + 273.15 = 285.15 K


Formula: V1/T1 = V2/T2


Solution:

Since the pressure and amount of gas remain unchanged, Charles's law can be used.

Charles's law can be expressed as: V1 / T1 = V2 / T2

V1T2 = V2T1

To find the resulting volume of the balloon, solve the equation for V2:

V2 = V1T2 / T1

V2 = (2.25 L × 285.15 K) / (313.15 K) = 2.05 L

V2 = 2.05 L


Answer: The volume of the balloon will be 2.05 L